Circular No. 6498

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996bp IN MCG –06-04-053 R. H. McNaught, Anglo-Australian Observatory, reports his discovery,

on an R plate exposed by F. G. Watson with the U.K. Schmidt on Aug. 11.8 UT, of an apparent supernova (mag ∼ 18) located at α = 1^{h}35^{m}20^{s}.80, δ = −32^{o}46′51′′.0 (equinox 2000.0; poorly placed near the edge of the plate, uncertainty 0′′.6 in each coordinate), which is 8′′.5 west and 15′′.4 south of the elliptical galaxy’s center. SN 1996bp is also present on a weak U film taken one day earlier (Aug. 10.8), with the supernova being much more prominent in relation to nearby stars. A nearby star (R ∼ 17, but fainter than SN 1996bp in U) has position end figures 22^{s}.51, 47′28′′.6. The supernova is not present on the ESO R or B surveys or on several U.K. Schmidt exposures, the latest being an R survey plate taken on 1991 Nov. 27.

SUPERNOVA 1996bo IN NGC 673 T. Hirayama, National Astronomical Observatory, Tokyo, reports that

K. Okazaki, Kahoku-machi, Yamagata-ken, independently discovered this object (cf. IAUC 6497) at mag ∼ 15 on an unfiltered CCD image taken with a 0.28-m reflector on Oct. 24.65 UT.

COMET C/1996 N1 (BREWINGTON)

Ephemeris extension to IAUC 6466 (elements MPC 27882):

1996TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Nov. 3 21^{h}06.^{m}50 +47^{o}47.^{′}9 1.216 1.775 106.^{o}6 32.^{o}4 13.2
8 21 28.18 +45 18.5 1.277 1.838 107.7 30.9 13.5
13 21 47.57 +42 51.9 1.344 1.902 108.2 29.6 13.8
18 22 05.00 +40 31.8 1.420 1.965 108.0 28.6 14.2
23 22 20.77 +38 20.7 1.502 2.028 107.3 27.7 14.5
28 22 35.16 +36 19.9 1.592 2.091 106.0 27.0 14.8
Dec. 3 22 48.39 +34 30.2 1.687 2.153 104.3 26.3 15.1
8 23 00.66 +32 51.6 1.788 2.216 102.2 25.8 15.4
13 23 12.12 +31 23.9 1.894 2.278  99.8 25.2 15.8
18 23 22.91 +30 06.6 2.004 2.340  97.2 24.7 16.0
23 23 33.12 +28 58.9 2.118 2.402  94.4 24.1 16.3

Visual m_{1} estimates: Oct. 9.14 UT, 12.2 (R. Keen, Mt. Thorodin, CO, 0.32-m reflector); 13.79, 11.9 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 20.04, 13.8: (J. Carvajal, Teruel, Spain, 0.44-m reflector).

1996 October 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green